OpenStudy (afifarahman):

The sides of a right angled triangle are x cm, (x+1)cm and (x+3)cm. Show that x^2 - 4x -8 = 0 and solve this equation to find the lengths of the sides of the triangle. give your answers correct to two decimal places.

OpenStudy (raden):

as the longest side is (x+3), right ? then (x+3) as the hypotenuse. use hte pythagorean theorem : x^2 + (x+1)^2 = (x+3)^2 simplify then solve for x

OpenStudy (afifarahman):

Thanks !

OpenStudy (raden):

btw, x^2 - 4x -8 = 0cant be factored. so you can use quadratic formula to solve for x. x = (-b +-sqrt(b^2 - 4ac))/(2a) here a = a, b = -4 and c = -8 you're welcome

OpenStudy (raden):

damn typo, i meant a = 1